# Estimating spectral radius of Bell-type operator via finite dimensional approximation of orthogonal projections

### v1

@@ Abstract
In quantum mechanics, the following problem is important: For two orthogonal projections $P$ and $Q$ on a separable infinite dimensional Hilbert space $V$, estimate the spectral radius $\rho([A,B])$, where $A=2P-\operatorname{id}_{V},B=2Q-\operatorname{id}_{V}$. This problem is known to be equivalent to determining the spectral radius of the Bell-CHSH operator.
To approach this problem, we give a method to approximate orthogonal projections on $V$ by those on finite dimensional subspaces.
This method provides an upper bound and a lower bound for $\rho([A,B])$, which become exact when the matrix representations of $P$ and $Q$ are “constant-angle one-shifted form”.
@@ 1  Introduction
From these observations, we conclude that if $V$ is finite dimensional, let $\lambda$ be the eigenvalue of $A+B$ whose absolute value is closest to $\sqrt{2}$, then the following holds:
@@
Second, for two orthogonal projections $P$ and $Q$, we provide an explicit formula for a finite dimensional approximation of the pair $(P,Q)$. This can be summarized as follows. We prove the existence of an orthogonal direct sum decomposition of $V$ into invariant subspaces under $P$ and $Q$, in which every infinite dimensional subspace admits an explicit finite dimensional approximation of the pair $(P,Q)$ as described in Theorem 1.1. We refer to the infinite dimensional matrix representations admitting this approximation as the one-shifted form. For the definition of one-shifted form, see Definition 4.1. From the viewpoint of group representation theory, two orthogonal projections having a one-shifted form yield a cyclic unitary representation of $D_{\infty}$. Theorem 1.1 thus provides an explicit formula that decomposes every infinite dimensional unitary representation of $D_{\infty}$ into cyclic representations. In other words, the one-shifted form (Definition 4.1) extends Jordan’s lemma to the infinite dimensional case. For details of this theorem, see Section 4.
Theorem 1.1 (Approximation of one-shifted form).
There exists an orthonormal basis $\mathcal{U}=\{u_{i}\}_{i=1}^{\infty}$ and a pair of sequences $\Theta=\{\theta_{i}\}_{i=1}^{\infty},\Omega=\{\omega_{i}\}_{i=1}^{\infty}\subset(0,\pi)$ such that, for any vector $v\in V$,
@@
Third, when $P$ and $Q$ admit a one-shifted form, we derive the inequalities as stated in Theorem 1.2. This theorem estimates the spectral radius of $[A,B]$ from above and below using the eigenvalues of symmetric tridiagonal matrices approximating $A+B$, based on the explicit formula given in Theorem 1.1. Theorem 1.2 is an extension of equation (1.1) to infinite dimensional Hilbert spaces.
@@
where for $n\in\mathbb{N}$, we define $A_{n}=2P_{n}-\operatorname{id}_{V_{n}},B_{n}=2Q_{n}-\operatorname{id}_{V_{n}}$, and let $\lambda_{n}$ be the eigenvalue of $A_{n}+B_{n}$ whose absolute value is closest to $\sqrt{2}$.
@@ 2.1 Notations
    $\tau=\tau_{P_{1},\dots,P_{m}}:$ The unitary representation defined by Fact 2.1, for orthogonal projections $P_{1},\dots,P_{m}$ on a Hilbert space $V$.
@@ 2.2 Unitary representations of $\mathbb{Z}_{2}*\cdots*\mathbb{Z}_{2}$
The following Fact 2.2 is well known concerning the irreducible decomposition of finite dimensional unitary representations of $D_{\infty}$. In quantum information theory, Fact 2.2 is referred to as Jordan’s lemma.
Fact 2.2 (Jordan’s lemma).
@@
The following result is a consequence of Fact 2.2 and elementary calculations.
Fact 2.3.
@@
    Let $\lambda$ be the eigenvalue of $A+B$ whose absolute value is closest to $\sqrt{2}$. Then
@@ 2.4.2 Explicit eigenvalues and eigenvectors of Toeplitz matrices
Yueh [16] studied the explicit formula for eigenvalues of the following $n\times n$ tridiagonal matrix:
@@
where $n\in\mathbb{N}$ and $a,b,\gamma,\delta,\alpha,\beta\in\mathbb{C}$. Define $\mu$ by $\mu=\sqrt{\dfrac{a}{c}}$. According to Theorem 3.1 of [16], for any eigenvalue $\lambda$ of $X$ there exists $\phi\in[0,2\pi)$ such that
$\lambda=\left\{\begin{array}[]{ll}b+2c\mu\cos\phi&\quad\sin\phi\neq 0,\\ b\pm 2c\mu&\quad\sin\phi=0,\end{array}\right$.
and if $\sin\phi\neq 0$ then $\phi$ satisfies the following equation:
@@
Moreover, according to (3.3) of [16], for any eigenvalue $\lambda$ of $X$ such that $\sin\phi\neq 0$, and for any eigenvector $u$ of $X$ corresponding to $\lambda$,
@@ 3.3 Application: New proof of Tsirelson’s inequality
If $\dim V_{1}=1$, (2.2) follows by simple calculus. Thus, we can assume that $\dim V_{1}=2$ and $\tau_{P_{1,1},P_{1,2}}$ is irreducible. From Fact 2.2 and Fact 2.3, there exists $x\in(0,1)$ such that
@@
From Fact 2.3, the eigenvalues of $C$ are $2\sqrt{x},2\sqrt{x}$ and the eigenvalues of $D$ are $2\sqrt{1-x},2\sqrt{1-x}$. Therefore, the spectral radius $\rho(\mathcal{B})$ is at most $2\sqrt{x}+2\sqrt{(1-x)}$. From simple differential calculus, we conclude that
@@
From Example 3.8 and Corollary 3.10, we may assume that $\dim V_{1}<\infty$. Moreover, from Fact 2.2, there exists a finite family $\{W_{i}\}_{i=1}^{m}$ of invariant subspaces under $P_{1,1}$ and $P_{1,2}$ such that $\dim W_{i}\leq 2$ for any $i\in\{1,\dots,m\}$ and $V_{1}=\bigoplus_{i=1}^{m}W_{i}$ and
@@
Tsirelson [13] established an ingenious technique using the Clifford algebra to reduce Tsirelson’s inequality to the case where $\dim V_{i}\leq 2$ $(i=1,2)$ in Theorem 2.4. Our proof presents another approach that reduces Tsirelson’s inequality to the two dimensional case, using Corollary 3.10 and Fact $2.$
@@ 4.2 Extension of Jordan’s lemma
It is straightforward that $\mathcal{T}$ is not empty. We assume that the order of $\mathcal{T}$ is defined by set inclusion. Let $T$ be a totally ordered subset of $\mathcal{T}$. It follows directly that $\bigcup_{D\in T}D\in\mathcal{T}$. From Zorn’s Lemma, the set $\mathcal{T}$ has a maximal element $D$. We define $V_{0}$ by $V_{0}=\bigoplus_{W\in D}W.$ To derive a contradiction, we assume $V_{0}\neq V$. If $Q|V_{0}^{\perp}=0$ or $V_{0}^{\perp}$ admits a nonzero finite dimensional invariant subspace of $V$ under $\tau_{P,Q}$, from Fact 2.2, the subspace $V_{0}^{\perp}$ has an at most two dimensional invariant subspace $W$ under $\tau_{P,Q}$. Then $W\in\mathcal{B}$, and hence $D\cup\{W\}\in\mathcal{T}$. That contradicts the fact that $D$ is a maximal element of $\mathcal{T}$.
@@
Let $W$ be a nonzero invariant subspace under $\tau_{P_{1},P_{2}}$. From Fact 2.2, it is enough to show that $\dim W\geq 3$. Since $W$ is nonzero, there exists $w\in W\setminus\{0\}$. From statement (ii), there are $m\in\mathbb{N}$ and a sequence $i_{1},\dots,i_{m}\in\{1,2\}$ and $\alpha\in\mathbb{C}$ such that the first component of $\alpha A_{i_{1}}\dots A_{i_{m}}w$ is $1$. Without loss of generality, we may assume that $w_{1}=1$. Since $w\in\ell^{2}$, we have $\lim_{n\to\infty}w_{n}=0$. Therefore, there exists $i_{0}\in[2,\infty)\cap\mathbb{N}$ such that $|w_{i_{0}}|\neq 1$ and $w_{j}=1$ for any $j\in\{1,\dots,i_{0}-1\}$. From statement (ii), there exists a finite sequence $j_{1},\dots,j_{n}\in\{1,2\}$ such that $A_{j_{1}}\dots A_{j_{n}}w=(1,w^{\prime}_{2},\dots),A_{1}A_{j_{1}}\dots A_{j_{n}}w=(w^{\prime}_{2},1,\dots),$ where $|w^{\prime}_{2}|\neq 1$. Since $(w^{\prime}_{2})^{2}\neq 1$, these two vectors are linearly independent. Hence, there exist $u,v\in W$ such that $u=(1,0,u_{3},\dots),v=(0,1,v_{3},\dots)$.
@@ 5.1 Upper and lower estimates of the spectral radius
In Theorem 5.3, which is the main result of this section, we provide upper and lower bounds for $\rho([A,B])$. The upper bound is established in the general setting, without assuming Case 1 or Case 2. The lower bound is derived under the assumption that either Case 1 or Case 2 holds. As a preparation for Theorem 5.2, we first identify $F(A+B)$ in Case 1. In Case 1, Fact 2.3 allows us to directly identify set $F(A+B)$, as follows.
Lemma 5.2.
@@
    $\{I_{n}\}_{n=1}^{\infty}$ be a family of finite subsets of $I$ such that $I_{n}\subset I_{n+1}$ for any $n\in\mathbb{N}$ and $\bigcup_{n=1}^{\infty}I_{n}=I$,
  3. (S5)
    $\mathcal{A}$ be a finite dimensional approximation of $(V,\{P,Q\})$ defined by
    $\mathcal{A}=\{(V_{n},\{P_{n},Q_{n}\})\}_{n=1}^{\infty},V_{n}=\bigoplus_{i\in I_{n}}W_{i},P_{n}=\bigoplus_{i\in I_{n}}P|W_{i},Q_{n}=\bigoplus_{i\in I_{n}}Q|W_{i},$
then we have
We give upper and lower estimates of the spectral radius of the commutator $[A,B]$, as follows.
@@
    Let $\lambda_{n}$ be the eigenvalue of $A_{n}+B_{n}$ whose absolute value is closest to $\sqrt{2}$ for $n\in\mathbb{N}$. Then
@@
    Assume that $(V,\{P,Q\})$ satisfies the conditions (S1)-(S3), (A1) in Lemma 5.2. Then
@@
Proof of (i):.
Let $\{(V_{\phi(j)},\{P_{\phi(j)},Q_{\phi(j)}\})\}_{j=1}^{\infty}$ be a subsequence of $\mathcal{A}$ such that $\{b(\lambda_{\phi(j)})\}_{j=1}^{\infty}$ converges, and let $v\in V$ and $\varepsilon>0$. From Proposition 3.4, there exists $n_{0}\in\mathbb{N}$ such that for any $n\in[n_{0},\infty)\cap\mathbb{N}$, we have $|([A,B]v,v)|-\varepsilon<|([A_{n},B_{n}]\pi_{n}v,\pi_{n}v)|$. From Fact 2.3, for any $n\in[n_{0},\infty)\cap\mathbb{N}$
@@
Let $i\in\mathbb{N}$. Then all subdiagonal elements of $A_{i}+B_{i}$ are nonzero. From Fact 2.5, the eigenvalues of $A_{i}+B_{i}$ are distinct. From Fact 2.2 and Fact 2.3, there exists a two dimensional irreducible invariant subspace $W$ under $\tau_{P_{i},Q_{i}}$ such that $(A_{i}+B_{i})|W$ has eigenvalues $\lambda_{i},\lambda_{i}$ and $u_{i},v_{i}\in W$. From Fact 2.3, there exist $\alpha_{i},\beta_{i}\in\mathbb{C}$ such that $|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1$ and
@@
where $w_{i}=\alpha_{i}u_{i}+\beta_{i}v_{i}$. From the definition of the spectral radius of $[A,B]$, we have $|([A,B]w_{i},w_{i})|\leq\rho([A,B])$. Let $\varepsilon>0$. For any sufficiently large $i\in\mathbb{N}$, the inequality $|[A,B]w_{i}-[A_{i},B_{i}]w_{i}|<\varepsilon$ holds. Then

### v2

# Estimating the spectral radius of Bell-type operator via finite dimensional approximation of orthogonal projections

@@ Abstract
We establish a new decomposition formula for two orthogonal projections $P$ and $Q$ on a separable Hilbert space $V$. This formula yields an orthogonal direct sum decomposition of $V$ into invariant subspaces under $P$ and $Q$, each of which is either at most two dimensional or infinite dimensional. On every infinite dimensional component, the pair $(P,Q)$ admits a matrix representation that we call the “one-shifted form”. This representation diagonalizes both $P$ and $Q$ into blocks of size at most two, and moreover, both projections can be explicitly approximated by orthogonal projections on finite dimensional subspaces. This approximation scheme offers a way to derive infinite dimensional results from their finite dimensional counterparts and is also useful in numerical computations.
This decomposition provides a useful framework for analyzing a wide range of problems involving two orthogonal projections in infinite dimensions. In particular, several spectral problems for operators generated by $P$ and $Q$ (including polynomials in $P$ and $Q$) can be reduced to the case where the pair admits a one-shifted form.
More concretely, we can estimate the spectral radius $\rho([P,Q])$, which is equivalent to estimating the spectral radius of the Bell-CHSH operator, a quantity of fundamental importance in quantum mechanics. We provide an upper bound and a lower bound for $\rho([P,Q])$, which become exact when the matrix representations of $P$ and $Q$ are in “constant-angle one-shifted form”.
@@ 1  Introduction
From these observations, we obtain the following proposition.
Proposition 1.1.
Assume that $V$ is finite dimensional and let $\lambda$ be an eigenvalue of $A+B$ whose square is closest to $2$. Then
@@
Second, for two orthogonal projections $P$ and $Q$, we provide an explicit formula for a finite dimensional approximation of the pair $(P,Q)$. This can be summarized as Theorem 1.2. For details of this theorem, see Section 4.
Theorem 1.2.
There exists an orthogonal direct sum decomposition of $V$ into invariant subspaces under $P$ and $Q$, each of which is either at most two dimensional or infinite dimensional. Every such infinite dimensional component admits an explicit finite dimensional approximation of the pair $(P,Q)$ in the following sense. There exists an orthonormal basis $\mathcal{U}=\{u_{i}\}_{i=1}^{\infty}$ and a pair of sequences $\Theta=\{\theta_{i}\}_{i=1}^{\infty},\Omega=\{\omega_{i}\}_{i=1}^{\infty}\subset(0,\pi)$ such that, for any vector $v\in V$,
@@
We refer to the infinite dimensional matrix representations admitting this approximation as the one-shifted form. For the definition of one-shifted form, see Definition 4.1.
Theorem 1.2 is the main result of this paper. From the viewpoint of group representation theory, two orthogonal projections having a one-shifted form yield a cyclic unitary representation of $D_{\infty}$. Theorem 1.2 thus provides an explicit formula that decomposes every infinite dimensional unitary representation of $D_{\infty}$ into cyclic representations. In other words, the one-shifted form (Definition 4.1) extends Jordan’s lemma to the infinite dimensional case. As a consequence, a variety of problems concerning infinite dimensional unitary representations of $D_{\infty}$ (two orthogonal projections in an infinite dimensional Hilbert space)—such as studying the spectrum of operators given by polynomials in $P$ and QQ—can be reduced to the case where the pair admits a one-shifted form.
@@
Third, when $P$ and $Q$ admit a one-shifted form, we derive the inequalities as stated in Theorem 1.3. This theorem estimates the spectral radius of $[A,B]$ from above and below using the eigenvalues of symmetric tridiagonal matrices approximating $A+B$, based on Theorem 1.2. Theorem 1.3 is an extension of equation (1.1) to infinite dimensional Hilbert spaces.
@@
where for $n\in\mathbb{N}$, we define $A_{n}=2P_{n}-\operatorname{id}_{V_{n}},B_{n}=2Q_{n}-\operatorname{id}_{V_{n}}$, and let $\lambda_{n}$ be an eigenvalue of $A_{n}+B_{n}$ such that $\lambda_{n}^{2}$ is closest to $2$.
@@
After the submission of the first version of this preprint [14], Böttcher and Spitkovsky [15] obtained an infinite dimensional extension of Proposition 1.1, from which a part of our Theorem 1.3 can be alternatively derived. Furthermore, also in [15], they kindly pointed out minor errors in the earlier version of our preprint and helped us correct them, for which we are sincerely grateful.
@@ 2.1 Notations
    $\tau=\tau_{P_{1},\dots,P_{m}}:$ The unitary representation defined by Proposition 2.1, for orthogonal projections $P_{1},\dots,P_{m}$ on a Hilbert space $V$.
@@ 2.2 Unitary representations of $\mathbb{Z}_{2}*\cdots*\mathbb{Z}_{2}$
The following Proposition 2.2 is well known concerning the irreducible decomposition of finite dimensional unitary representations of $D_{\infty}$. In quantum information theory, Proposition 2.2 is referred to as Jordan’s lemma.
Proposition 2.2 (Jordan’s lemma).
@@
The following result is a consequence of Proposition 2.2 and elementary calculations.
Proposition 2.3.
@@
    Let $\lambda$ be an eigenvalue of $A+B$ such that $\lambda^{2}$ is closest to $2$. Then
@@ 2.4.2 Explicit eigenvalues and eigenvectors of Toeplitz matrices
Yueh and Cheng [18] provides a method for constructing eigenpairs of the following $n\times n$ tridiagonal matrix:
@@
where $n\in\mathbb{N}$ and $a,b,\gamma,\delta,\alpha,\beta\in\mathbb{C}$. We focus on the case when $a=c\in\mathbb{R}$ and $b\in\mathbb{R}$. Define $\mu$ by $\mu=\sqrt{\dfrac{a}{c}}$. Assume that we find $\phi\in[0,2\pi)$ such that $\sin\phi\neq 0$ and $\phi$ satisfies the following equation:
@@
Then according to Theorem 3.1 of [18],
$\lambda=b+2c\mu\cos\phi$
is an eigenvalue $\lambda$ of $X$. Moreover, from (3.3) of [18], for any eigenvector $u$ of $X$ corresponding to $\lambda$,
@@ 3.3 Application: New proof of Tsirelson’s inequality
If $\dim V_{1}=1$, (2.2) follows by simple calculus. Thus, we can assume that $\dim V_{1}=2$ and $\tau_{P_{1,1},P_{1,2}}$ is irreducible. From Proposition 2.2 and Proposition 2.3, there exists $x\in(0,1)$ such that
@@
From Proposition 2.3, the eigenvalues of $C$ are $2\sqrt{x},2\sqrt{x}$ and the eigenvalues of $D$ are $2\sqrt{1-x},2\sqrt{1-x}$. Therefore, the spectral radius $\rho(\mathcal{B})$ is at most $2\sqrt{x}+2\sqrt{(1-x)}$. From simple differential calculus, we conclude that
@@
From Example 3.8 and Corollary 3.10, we may assume that $\dim V_{1}<\infty$. Moreover, from Proposition 2.2, there exists a finite family $\{W_{i}\}_{i=1}^{m}$ of invariant subspaces under $P_{1,1}$ and $P_{1,2}$ such that $\dim W_{i}\leq 2$ for any $i\in\{1,\dots,m\}$ and $V_{1}=\bigoplus_{i=1}^{m}W_{i}$ and
@@
Tsirelson [13] established an ingenious technique using the Clifford algebra to reduce Tsirelson’s inequality to the case where $\dim V_{i}\leq 2$ $(i=1,2)$ in Theorem 2.4. Our proof presents another approach that reduces Tsirelson’s inequality to the two dimensional case, using Corollary 3.10 and Proposition $2.$
@@ 4.2 Extension of Jordan’s lemma
It is straightforward that $\mathcal{T}$ is not empty. We assume that the order of $\mathcal{T}$ is defined by set inclusion. Let $T$ be a totally ordered subset of $\mathcal{T}$. It follows directly that $\bigcup_{D\in T}D\in\mathcal{T}$. From Zorn’s Lemma, the set $\mathcal{T}$ has a maximal element $D$. We define $V_{0}$ by $V_{0}=\bigoplus_{W\in D}W.$ To derive a contradiction, we assume $V_{0}\neq V$. If $Q|V_{0}^{\perp}=0$ or $V_{0}^{\perp}$ admits a nonzero finite dimensional invariant subspace of $V$ under $\tau_{P,Q}$, from Proposition 2.2, the subspace $V_{0}^{\perp}$ has an at most two dimensional invariant subspace $W$ under $\tau_{P,Q}$. Then $W\in\mathcal{B}$, and hence $D\cup\{W\}\in\mathcal{T}$. That contradicts the fact that $D$ is a maximal element of $\mathcal{T}$.
@@
Let $W$ be a nonzero invariant subspace under $\tau_{P_{1},P_{2}}$. From Proposition 2.2, it is enough to show that $\dim W\geq 3$. Since $W$ is nonzero, there exists $w\in W\setminus\{0\}$. From statement (ii), there are $m\in\mathbb{N}$ and a sequence $i_{1},\dots,i_{m}\in\{1,2\}$ and $\alpha\in\mathbb{C}$ such that the first component of $\alpha A_{i_{1}}\dots A_{i_{m}}w$ is $1$. Without loss of generality, we may assume that $w_{1}=1$. Since $w\in\ell^{2}$, we have $\lim_{n\to\infty}w_{n}=0$. Therefore, there exists $i_{0}\in[2,\infty)\cap\mathbb{N}$ such that $|w_{i_{0}}|\neq 1$ and $w_{j}=1$ for any $j\in\{1,\dots,i_{0}-1\}$. From statement (ii), there exists a finite sequence $j_{1},\dots,j_{n}\in\{1,2\}$ such that $A_{j_{1}}\dots A_{j_{n}}w=(1,w^{\prime}_{2},\dots),A_{1}A_{j_{1}}\dots A_{j_{n}}w=(w^{\prime}_{2},1,\dots),$ where $|w^{\prime}_{2}|\neq 1$. Since $(w^{\prime}_{2})^{2}\neq 1$, these two vectors are linearly independent. Hence, there exist $u,v\in W$ such that $u=(1,0,u_{3},\dots),v=(0,1,v_{3},\dots)$.
@@ 5.1 Upper and lower estimates of the spectral radius
Notation 5.2.
@@
    $\{I_{n}\}_{n=1}^{\infty}$ be a family of finite subsets of $I$ such that $I_{n}\subset I_{n+1}$ for any $n\in\mathbb{N}$ and $\bigcup_{n=1}^{\infty}I_{n}=I$.
Under these assumption we define $\mathcal{A}$ by
$\mathcal{A}$ $=\{(V_{n},\{P_{n},Q_{n}\})\}_{n=1}^{\infty},$
$V_{n}$ $=\bigoplus_{i\in I_{n}}W_{i},P_{n}=\bigoplus_{i\in I_{n}}P|W_{i},Q_{n}=\bigoplus_{i\in I_{n}}Q|W_{i}\quad(n=1,2,\dots)$
Remak that $\mathcal{A}$ is a finite dimensional approximation of $(V,\{P,Q\})$.
In Theorem 5.3, which is the main result of this section, we provide upper and lower bounds for $\rho([A,B])$. The upper bound is established in the general setting, without assuming Case 1 or Case 2. The lower bound is derived under the assumption that either Case 1 or Case 2 holds.
@@
    Let $\lambda_{n}$ be an eigenvalue of $A_{n}+B_{n}$ such that $\lambda_{n}^{2}$ is closest to $2$ for $n\in\mathbb{N}$. Then
@@
    Assume that $(V,\{P,Q\})$ satisfies the conditions (S1)-(S3), (A1) in Notation 5.2 and $\mathcal{A}$ is the finite dimensional approximation of $(V,\{P,Q\})$ defined in Notation 5.2. Then
@@
After the submission of the first version of this preprint [14], Böttcher and Spitkovsky [15] obtained an infinite dimensional extension of Proposition 1.1. Hereafter, we provide a proof of Theorem 5.3 without their extension.
As a preparation for the proof of Theorem 5.3, we first identify $F(A+B)$ in Case 1. In Case 1, Proposition 2.3 allows us to directly identify set $F(A+B)$, as follows.
Lemma 5.1.
Assume the setups in Notation 5.2. Then we have
We give the proof of Theorem 5.3 as follows.
Proof of (i) of Theorem 5.3:.
Let $\{(V_{\phi(j)},\{P_{\phi(j)},Q_{\phi(j)}\})\}_{j=1}^{\infty}$ be a subsequence of $\mathcal{A}$ such that $\{b(\lambda_{\phi(j)})\}_{j=1}^{\infty}$ converges, and let $v\in V$ and $\varepsilon>0$. From Proposition 3.4, there exists $n_{0}\in\mathbb{N}$ such that for any $n\in[n_{0},\infty)\cap\mathbb{N}$, we have $|([A,B]v,v)|-\varepsilon<|([A_{n},B_{n}]\pi_{n}v,\pi_{n}v)|$. From Proposition 2.3, for any $n\in[n_{0},\infty)\cap\mathbb{N}$
@@
Let $i\in\mathbb{N}$. Then all subdiagonal elements of $A_{i}+B_{i}$ are nonzero. From Proposition 2.5, the eigenvalues of $A_{i}+B_{i}$ are distinct. From Proposition 2.2 and Proposition 2.3, there exists a two dimensional irreducible invariant subspace $W$ under $\tau_{P_{i},Q_{i}}$ such that $(A_{i}+B_{i})|W$ has eigenvalues $\lambda_{i},\lambda_{i}$ and $u_{i},v_{i}\in W$. From Proposition 2.3, there exist $\alpha_{i},\beta_{i}\in\mathbb{C}$ such that $|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1$ and
@@
where $w_{i}=\alpha_{i}u_{i}+\beta_{i}v_{i}$. From the definition of the spectral radius of $[A,B]$, we have $|([A,B]w_{i},w_{i})|\leq\rho([A,B])$. Let $\varepsilon>0$.
We show that, for any sufficiently large $i\in\mathbb{N}$, the inequality $|[A,B]w_{i}-[A_{i},B_{i}]w_{i}|<\varepsilon$ holds. From the definition of $\{u_{i}\}_{i=1}^{\infty}$ and $\{v_{i}\}_{i=1}^{\infty}$, there exists $i_{0}\in\mathbb{N}$ such that for any $i\in[i_{0},\infty)\cap\mathbb{N}$
$\max\{\|(A-A_{i})u_{i}\|,\|(B-B_{i})u_{i}\|,\|(A-A_{i})v_{i}\|,\|(B-B_{i})v_{i}\|\}<\dfrac{\varepsilon}{8}.$
Let $i\in[i_{0},\infty)\cap\mathbb{N}$. Since $|\alpha_{i}|,|\beta_{i}|\leq 1$, it follows that
$\|Bw_{i}-B_{i}w_{i}\|\leq|\alpha_{i}|\|(B-B_{i})u_{i}\|+|\beta_{i}|\|(B-B_{i})v_{i}\|<\dfrac{\varepsilon}{4}.$
Because $A$ is unitary, we obtain
$\|ABw_{i}-AB_{i}w_{i}\|<\dfrac{\varepsilon}{4}.$
Since $W$ is invariant under $B_{i}$ and $B_{i}$ is unitary, we have $B_{i}w_{i}\in W$ and $\|B_{i}w_{i}\|=1$. From Proposition 2.3, there exists $\alpha_{i}^{\prime},\beta_{i}^{\prime}\in\mathbb{C}$ such that
$B_{i}w_{i}=\alpha_{i}^{\prime}u_{i}+\beta_{i}^{\prime}v_{i},\quad|\alpha_{i}^{\prime}|^{2}+|\beta_{i}^{\prime}|^{2}=1.$
Since $|\alpha_{i}^{\prime}|,|\beta_{i}^{\prime}|\leq 1$, it follows that
$\|AB_{i}w_{i}-A_{i}B_{i}w_{i}\|\leq|\alpha_{i}^{\prime}|\|(A-A_{i})u_{i}\|+|\beta_{i}^{\prime}|\|(A-A_{i})v_{i}\|<\dfrac{\varepsilon}{4}.$
Therefore, we obtain $\|ABw_{i}-A_{i}B_{i}w_{i}\|<\dfrac{\varepsilon}{2}$. Following the same reasoning as above, we have $\|BAw_{i}-B_{i}A_{i}w_{i}\|<\dfrac{\varepsilon}{2}$. We conclude that $\|[A,B]w_{i}-[A_{i},B_{i}]w_{i}\|<\varepsilon$.
Thus we have